OpenStudy (anonymous):

RATIO & PROPORTIONS: A poster is 16 inches long and 24 inches wide. It is enlarged so that the shorter side is 24 inches. What is the length of the longer side?

OpenStudy (anonymous):

36 inches, I think.

OpenStudy (anonymous):

I have this :\[\frac{ 16 }{ 40 } = \frac{ x }{ 24 } x= 9.6\]

OpenStudy (anonymous):

where'd you get 40 from?

OpenStudy (anonymous):

the longer side is 36 inches

OpenStudy (anonymous):

yeah, the answer should be 36 inches.

OpenStudy (anonymous):

I have no clue, i copied it from my teacher when he was writing it on the board. i need help.

OpenStudy (anonymous):

This is what you were given right? 16/24 =2/3 24/x =2/3 In this equation, x needs to be 36 so it will also be 2/3 as a simplified fraction.

OpenStudy (anonymous):

Did you get it?